OpenStudy (anonymous):

solve for x 0=(24x^3)-(72x^2)-(186x)+360 solve for x 0=(24x^3)-(72x^2)-(186x)+360 @Mathematics

OpenStudy (anonymous):

x1= -5/2, x2=3/2, x3=4

OpenStudy (anonymous):

(24x^3)-(72x^2)-(186x)+360 = x^3 - 3x^2 - 31x/4 +15 First we find one root by guessing or plotting the polynomial, and we see that one root is x = 4. Then we solve the rest as follows: (x^3 - 3x^2 - 31x/4 +15)/(x-4) = x^2 + x - 15/4 (x-4)(x^2 + x - 15/4) = 0 d = b^2 -4ac = 1 - 4*1*(-15/4) = 16 x = (-b+-sqrt(d))/2a = (-1+-4)/2, so x = -5/2 or x = 3/2

OpenStudy (anonymous):

because we know that (x-4) = 0